OpenStudy (anonymous):

A model airplane is shot up from a platform 1 foot above the ground with an initial upward velocity of 56 feet per second. The height of the airplane above the ground after t seconds is given by the equation h=−16t²+56t+1, where h is the height of the airplane in feet and t is the time in seconds after it is launched. What is the maximum height the airplane reaches, to the nearest foot? A:30ft B:50ft C:60ft D:80ft

OpenStudy (whpalmer4):

\[h(t) = -16t^2+56t+1\] That's a parabola. The vertex of the parabola is the maximum height the airplane will reach.

OpenStudy (whpalmer4):

You can find the vertex of a parabola in the form \[y = ax^2+bx+c\] by evaluating \[x = -\frac{b}{2a}\]and then evaluating the function at that value of \(x\) to get the corresponding \(y\) value

OpenStudy (anonymous):

so x=-56/2(-16)

OpenStudy (whpalmer4):

good so far...

OpenStudy (anonymous):

x=7/4 is what i got from that

OpenStudy (whpalmer4):

yes, that's correct. what will the value of \(h(\frac{7}{4})=\)

OpenStudy (anonymous):

so we solved for t correct?

OpenStudy (whpalmer4):

\(t=7/4\), yes now you need to find the actual height at that time.

OpenStudy (anonymous):

\[h(\frac{ 7 }{ 4 }) = -16(\frac{ 7 }{ 4 }) + 56(\frac{ 7 }{ 4 }) +1\]

OpenStudy (whpalmer4):

no, the first term has t^2, right?

OpenStudy (anonymous):

wait -16(7/4) is squared

OpenStudy (whpalmer4):

snap :-)

OpenStudy (whpalmer4):

only the 7/4 is squared, not the -16

OpenStudy (anonymous):

ok so i got 28.5 would the anwser be A?

OpenStudy (anonymous):

cause it asks for the nearest foot

OpenStudy (anonymous):

i did this wrong... sorry to waste your time thanks for the help anyway

OpenStudy (whpalmer4):

Hmm, looks like you did the arithmetic incorrectly. \[h(\frac{7}{4}) = -16(\frac{7}{4})^2+56(\frac{7}{4})+1\]\[\qquad=-16(\frac{49}{16})+14*7 + 1\]\[\qquad=-49+98+1\]\[\qquad=50\] Here's a plot of the height vs. time, with grid lines added to show the exact vertex location.